OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0

OpenStudy (anonymous):

12x^2-8x+3x-2=0 12x^2-5x-2=0 x=[5+(25+96)^1/2]/24 x=[5+(119)^1/2]/24 and x=[5-(119)^1/2]/24

OpenStudy (campbell_st):

its simply a case of solving 4x + 1 = 0 and 3x - 2 = 0 this gives the solution to the quadratic

OpenStudy (anonymous):

this is a quadratic equation , and I have solved it using the formula for finding roots ax^2-bx+c=0 so x=[-(-b)+/-(b^2-4ac)]/2a What campbell and luis have posted are also completely correct , and infact are straightforward and shortcut method. So it depends what you want, straightforward method , or you have to solve it using the formula ( specifically for your tests or exams )

OpenStudy (anonymous):

thank you for your responses! yeah it's correct =D

OpenStudy (anonymous):

The question simply says "Solve"..it doesn't specifically state a method.

OpenStudy (anonymous):

Thats right uniden , but it depends , if this question comes in objective( choice ) paper then straightforward way is "d best" , but if it comes in some school test , and say its for 5 marks , then method indeed matters.

OpenStudy (anonymous):

your welcum fatima ! :-)